• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 19:46:21 on March 06, 2012

    Entry number 364251

    keyword=Hall Status

    Target people came out of the hall around 18:30. Requested sweep. Sweep in progress until around 19:15. Thereafter, going to beam permit. Presently ramping up target magnetic field to 2.5 T. The plan is to tune beam through the chicane once the target is ready.